### v1

# Synthesizing Efficient and Permissive Programmatic Runtime Shields for Neural Policies

Abstract.
With the increasing use of neural policies in control systems, ensuring their safety and reliability has become a critical software engineering task. One prevalent approach to ensuring the safety of neural policies is to deploy programmatic runtime shields alongside them to correct their unsafe commands. However, the programmatic runtime shields synthesized by existing methods are either computationally expensive or insufficiently permissive, resulting in high overhead and unnecessary interventions on the system. To address these challenges, we propose Aegis, a novel framework that synthesizes lightweight and permissive programmatic runtime shields for neural policies. Aegis achieves this by formulating the seeking of a runtime shield as a sketch-based program synthesis problem and proposing a novel method that leverages counterexample-guided inductive synthesis and Bayesian optimization to solve it. To evaluate Aegis and its synthesized shields, we use four representative control systems and compare Aegis with the current state-of-the-art. Our results show that the programmatic runtime shields synthesized by Aegis can correct all unsafe commands from neural policies, ensuring that the systems do not violate any desired safety properties at all times. Compared to the current state-of-the-art, Aegis’s shields exhibit a 2.1$\times$ reduction in time overhead and a 4.4$\times$ reduction in memory usage, suggesting that they are much more lightweight. Moreover, Aegis’s shields incur an average of 1.6$\times$ fewer interventions than other shields, showing better permissiveness.
Runtime Shielding, Deep Reinforcement Learning, Program Synthesis, Trustworthy Systems
copyright: acmlicensedjournal: TOSEMccs: Software and its engineering Formal software verificationccs: Computing methodologies Computational control theory
1. Introduction
Recent years have seen a notable increase in the use of deep neural networks (DNNs) to support control systems like autonomous vehicles (Song et al., 2022; Ahmed et al., 2021; Lopez et al., 2023). In these systems, DNNs, also referred to as neural policies (Zhu et al., 2019; Donti et al., 2021), process information from the physical environment and send commands to hardware components like actuators to regulate their behavior in accordance with desired specifications. Neural policies excel at handling complex control tasks, such as autonomous driving (Kiran et al., 2021; Muhammad et al., 2020), aerospace guidance (Chai et al., 2021; Dong et al., 2021), and even nuclear fusion (Degrave et al., 2022; Seo et al., 2024), often outperforming traditional control software (Song et al., 2022; Lillicrap et al., 2016; Duan et al., 2016; Rajeswaran et al., 2017). According to Microsoft’s projection, by 2025, 50% of industrial companies will embrace control systems with neural policies, increasing productivity by 20% and generating \$3.7 trillion in revenue (Microsoft, 2020).
While neural policies have proven effective in various domains, concerns about their safety have arisen: they can occasionally produce unexpected or inappropriate commands, causing system failures that violate desired safety specifications (Pang et al., 2022; Amodei et al., 2016; He et al., 2024; Pei et al., 2017; Song et al., 2022; Zhang et al., 2020). This risk is exemplified in safety-critical tasks, e.g. autonomous driving, where neural policies may fail to control vehicles at desired heading angles, resulting in scenarios such as vehicles veering off the road or even colliding with pedestrians. Such catastrophic accidents have been reported multiple times in real-world systems employing neural policies, such as Tesla Autopilot (Post, 2023; Journalism, 2023) and Waymo Driver (Verge, 2023; TechCrunch, 2023). In addition, certain neural policies in aerospace and manufacturing control systems have also demonstrated unsafe behavior (SEA, 2019; Telegraph, 2023), further highlighting the safety risks associated with using neural policies in safety-critical domains.
Figure 1. Overview of how a programmatic runtime shield works. The green line detects potentially unsafe commands from the neural policy, while the red lines correct them by returning safe ones instead.
Programmatic Runtime Shields.   To date, programmatic runtime shields have demonstrated effectiveness in ensuring the safety of neural policies (Zhu et al., 2019; Alshiekh et al., 2018; Xiong and Jagannathan, 2021; Guo et al., 2023). These shields are typically simple and verifiable programs that can run alongside the neural policy and promptly correct any inappropriate commands from the policy as it executes. Figure 1 illustrates how a programmatic runtime shield works within a control system using a neural policy. Specifically, the shield monitors commands from the neural policy and intervenes when it detects potentially unsafe commands that could cause the system to violate required safety properties (line 2, marked in green). Upon intervention, the shield corrects unsafe commands by replacing them with safe alternatives based on the current state of the physical environment (lines 3-4, marked in red). This ensures that the desired safety properties are maintained at all times.
Given that the synthesis of these shields is typically less costly than formal verification or exhaustive testing of neural policies, they prove to be practical in ensuring the safety of neural policies (Zhu et al., 2019; Könighofer et al., 2020; Xiong and Jagannathan, 2021). Additionally, the synthesis and operation of runtime shields usually require only the inputs and outputs of the policies, rather than the neural network parameters. This makes deploying runtime shields often the only viable option for correcting safety violations of neural policies, especially when retraining neural networks is prohibitively expensive or when accessing the parameters of the neural policies is restricted in the case of using proprietary products from third-party vendors.
What Makes Good Programmatic Runtime Shields?  Three critical attributes define the efficacy of programmatic runtime shields: effectiveness, efficiency, and permissiveness (Zhu et al., 2019; Bloem et al., 2015; Könighofer et al., 2017; Wu et al., 2019; Könighofer et al., 2020). First, a good shield should effectively correct any unsafe command from the neural policy, thereby ensuring the perpetuation of desired safety properties in the system. Second, as an additional component running alongside the neural policy, the shield should be efficient, minimizing time and memory overhead. Lastly, maintaining permissiveness is essential (Zhu et al., 2019; Könighofer et al., 2023, 2020). The shield’s judgment of whether a command from the neural policy is unsafe is not always accurate, and it may intervene unnecessarily when it encounters safe commands. The shield should be permissive, minimizing unnecessary intervention and allowing the neural policy to operate as freely as possible.
Deficiencies in any of these key attributes can have adverse consequences. For example, an inefficient runtime shield can consume excessive system resources, resulting in degraded system performance or crashes. This is particularly problematic in real-time systems (Kopetz and Steiner, 2022) like autonomous vehicles, where timely responses are critical, or in resource-constrained embedded systems (Roth et al., 2024; Panda et al., 1999) like drones, where memory is limited. In addition, a non-permissive shield can often interfere with safe commands, thereby degrading system stability (i.e., a system’s ability to achieve steady states in response to random inputs or disturbances, typically measured by required time steps (Zhu et al., 2019); see Section 5.1 for further discussion). Unfortunately, current state-of-the-art programmatic runtime shields still have limitations with respect to these key attributes, as we discuss next.
State-of-the-art and Its Limitations
Several methods for synthesizing programmatic runtime shields have emerged in recent years, among which Zhu et al.’s work (An Inductive Synthesis Framework for Verifiable Reinforcement Learning (Zhu et al., 2019), hereafter referred to as VRL) stands out as the state-of-the-art. VRL crafts programmatic runtime shields capable of effectively rectifying all unsafe commands to consistently uphold desired safety properties (see Section 2 for more details). However, these shields have limitations in the other two attributes. First, VRL’s shields involve operations with multiple high-degree polynomials, resulting in a relatively large time and memory overhead on the systems they operate within. Second, their lack of permissiveness often leads to unnecessary interventions in safe commands, consequently compromising system stability. These limitations highlight opportunities for improving VRL’s shields, necessitating our work to extend the state-of-the-art.
Our Solution: Aegis
We introduce Aegis, a novel method to synthesize efficient and permissive programmatic runtime shields for neural policies. The core idea of Aegis is to formulate the seeking of a shield as a sketch-based program synthesis problem and use counterexample-guided inductive synthesis (Solar-Lezama, 2008) and Bayesian optimization (Garnett, 2023) to solve it. These two techniques have been studied in program synthesis in the past (Abate et al., 2018; Češka et al., 2019; Saad et al., 2019). However, Aegis is the first to combine them to enhance the efficiency and permissiveness of runtime shields, offering a unique perspective to the field.
Concretely, Aegis begins with constructing a program sketch that implements a linear policy and a switching strategy. The linear policy is a linear function whose coefficients are parameters to be synthesized for ensuring the generation of safe commands. Meanwhile, the switching strategy formulates a linear inequality based on the disparity between the commands from the neural policy and the linear policy, along with a threshold (also a parameter to be synthesized), to determine when to invoke the linear policy for correcting unsafe commands. With the program sketch, Aegis uses a novel counterexample-guided inductive synthesis algorithm and a Bayesian optimization algorithm to derive the necessary parameters for the linear policy and the switching strategy, respectively. Once all the parameters are obtained, Aegis integrates them into the sketch, forming a programmatic runtime shield that runs alongside the neural policy to ensure safety.
Two thrusts make Aegis superior to the state-of-the-art. First, the shields synthesized by Aegis involve only linear functions instead of the high-degree polynomials in VRL’s shields. This design choice reduces computational complexity and overhead, making our shields much more efficient. Second, our Bayesian optimization algorithm assigns the optimal threshold to the switching strategy by minimizing unnecessary interventions, thus making our shields more permissive than those synthesized by VRL.
Evaluating Aegis
We evaluate Aegis primarily on the three key attributes of programmatic runtime shields: effectiveness, efficiency, and permissiveness. In addition, we also evaluate Aegis’s scalability, showing its ability to efficiently synthesize shields for neural policies of different sizes. Our evaluation uses 4 representative control systems and compares Aegis to VRL.
Our results show that the programmatic runtime shields synthesized by Aegis effectively correct all unsafe commands from neural policies, ensuring that the systems do not violate the desired safety properties at any time. Notably, compared to VRL, Aegis’s shields exhibit a 2.1$\times$ reduction in time overhead and a 4.4$\times$ reduction in memory overhead, suggesting they are much more efficient. Aegis’s shields also show improved permissiveness, incurring an average of 1.6$\times$ fewer interventions than those synthesized by VRL. We also gauge the ratio of necessary interventions to all interventions and find that, on average, our shields intervene when necessary 9.9% more frequently than VRL’s. In addition, Aegis shows superior scalability relative to VRL, reducing synthesis time by an average of 77.3%. Additionally, we provide an analysis of replacing neural policies with Aegis’s shields, confirming the necessity for the coexistence of neural policies and shields to maintain the system’s safe and stable operation.
Main Contributions
Our main contributions are:
  • We improve the efficiency as well as permissiveness of programmatic runtime shields for neural policies, addressing the limitations of the state-of-the-art.
  • We propose and implement Aegis, a novel method that synthesizes programmatic runtime shields via a sketch-based program synthesis approach.
  • We evaluate Aegis on four representative systems and show its superiority over the state-of-the-art, with significant improvements in all metrics.
Paper Structure
The rest of the paper is organized as follows. Section 2 covers the preliminary information of our work, and Section 3 describes in detail how Aegis works. Section 4 presents the experimental setup and analyzes the experimental results. Section 5 provides additional discussions about our work, as well as threats to validity. Section 6 presents related work. Section 7 concludes our paper, and presents our future work and replication package.
2. Preliminary
This section provides a brief introduction to neural policies, as well as the synthesis of programmatic runtime shields.
Neural Policy
Recent advances in neural networks have spurred their adoption in many control systems (Song et al., 2022; Ahmed et al., 2021; Lopez et al., 2022). These neural networks generate commands to control systems that interact with the physical environment to fulfill desired specifications, often referred to as neural policies (Zhu et al., 2019). For clarity, below we formally describe neural policies.
We begin by defining the physical environment as $\mathcal{E}=\langle\mathcal{S},\mathcal{S}_{0},\mathcal{S}_{u},\mathcal{C},f,r\rangle$. Here $\mathcal{S}$ is the set of all possible states of the environment, with $s\in\mathcal{S}$ denoting a single state. The subset $\mathcal{S}_{0}\subset\mathcal{S}$ specifies the initial states from which systems working in this environment can commence. $\mathcal{S}_{u}\subset\mathcal{S}$ is a set of unsafe states that systems must avoid, thereby defining the desired safety specifications. $\mathcal{C}$ represents a set containing all feasible control commands that neural policies can issue, where $c\in\mathcal{C}$ denotes an individual command. The environment dynamics function $f$ governs the transition from one environment state to another upon receiving a command from a policy. Additionally, the reward function $r$ maps a state and a command to a real value as a feedback signal to guide the training process of a neural policy.
Modern control systems are typically implemented with digital rather than analog techniques (Lee, 2015; Zhu et al., 2019), where signals are processed discretely over fixed time intervals. Thus, the state transition of an environment can be formulated with discrete and fixed time step $t$ using Euler’s method (Hildebrand, 1987; Zhu et al., 2019). Namely, a state $s_{t}$ transitions to the subsequent state $s_{t+1}$ after time $t$ by the following equation:
(1) $s_{t+1}=s_{t}+f(s_{t},c_{t})\times\Delta t$
Here, $c_{t}$ denotes a command generated by a neural policy $\pi$, which is implemented as a neural network that takes an environment state $s_{t}$ as input, i.e., $c_{t}=\pi(s_{t})$, and $\Delta t$ is the fixed time step. Typically, neural policies are trained using Deep Reinforcement Learning (DRL) algorithms (Li, 2017; Lillicrap et al., 2016), which optimize the policy’s parameters to maximize the expected discounted cumulative reward through interaction with the environment. Specifically, in an environment $\mathcal{E}$, a system begins at an initial state $s_{0}\in\mathcal{S}_{0}$, and its neural policy produces commands to control the system to interact with the environment until it reaches a termination state, $s_{T}$, after $T$ time steps. The interaction history is called a trajectory:
(2) $\tau:(\langle s_{0},c_{0}\rangle,\cdots,\langle s_{t},c_{t}\rangle,\langle s_{t+1},c_{t+1}\rangle\cdots,\langle s_{T},c_{T}\rangle)$
The discounted cumulative reward is the sum of the rewards multiplied by a discount factor $\gamma\in[0,1]$ over a trajectory, denoted as $R(\tau)=\sum_{i=0}^{T}r(s_{i},c_{i})\cdot\gamma^{i}$. The training objective of DRL is to find the optimal neural policy $\pi^{\ast}$ that maximizes the expected discounted cumulative reward over all trajectories, which can be formalized as follows:
(3) $\pi^{\ast}=\arg\max_{\pi}\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\right]$
The specific DRL algorithms for solving the above equation vary with different tasks and are beyond our paper’s scope. We refer interested readers to (Li, 2017; Lillicrap et al., 2016) for more details.
1def main_loop():
2 $\eta$, $\omega$ = read_environment()
[…]
5 motor.update($c$)
Figure 2. Main control loop in a quadcopter environment (simplified for readability). The green and red colors highlight the neural policy and runtime shield, respectively.
After training, the neural policy $\pi$ works to control the system on which it was trained. Figure 2 shows an example of the main control loop of a quadcopter using a neural policy. The control loop starts by retrieving the current state, i.e., the current attitude angle $\eta$ and angular velocity $\omega$ (line 2). Next, the state is fed into the neural policy to obtain the resulting commands (line 3, marked in green). The command is then transmitted to the quadcopter’s motor to adjust its attitude and angular velocity (line 5). This loop continuously iterates to drive the quadcopter.
Programmatic Runtime Shields and VRL
While neural policies outperform traditional control methods in adapting control systems to more complex tasks (Lillicrap et al., 2016; Hunt et al., 1992; Duan et al., 2016), there remain potential risks that neural policies may output unsafe commands, resulting in system failures, such as misregulating motor power to cause quadcopters to crash. Therefore, a runtime shield becomes imperative to collaborate with the neural policy, ensuring the system’s safety when the neural policy fails to do so. Line 4 (marked in red) in Figure 2 shows the integration of a programmatic runtime shield into a control loop. The shield processes the neural policy’s command and the current state as input, generating a safe command as a replacement if the neural policy’s command is deemed unsafe.
VRL (Zhu et al., 2019) is a state-of-the-art approach for synthesizing programmatic runtime shields. Figure 3 shows an example of its shield for a neural policy within a quadcopter system. The shield contains a high-degree polynomial expression in line 3 (marked in green), which is a barrier certificate (Prajna and Jadbabaie, 2004; Gulwani and Tiwari, 2008; Kong et al., 2013), along with a linear policy in line 4 (marked in red) that has been verified to generate safe commands. The barrier certificate is a proof that if its result for a given state is non-positive, then the system is safe in that state; otherwise, the system is unsafe. When VRL’s shield works, it first uses a known mathematical model of the environment to predict the next state with the neural policy’s command (line 2), and then checks the safety of the next state using the barrier certificate (line 3). If the certificate yields a positive result, entering the next state steered by the neural policy’s command can inevitably make the system unsafe. In such cases, the shield returns the command of the linear policy that has a guarantee of safety instead (line 4). Conversely, if the certificate’s result is non-positive, the shield has the neural policy’s command returned (line 6).
1def programmatic_runtime_shield($c$, $\eta$, $\omega$):
2 $\hat{\eta}$, $\hat{\omega}$ = environment_model($c$, $\eta$, $\omega$)
[…]
5 else:
6 return $c$
Figure 3. A runtime shield synthesized by VRL for the neural policy in a quadcopter system. Line 3 (marked in green) is a barrier certificate that identifies unsafe states, while a linear policy in line 4 (marked in red) outputs safe commands.
1def programmatic_runtime_shield($c$, $\eta$, $\omega$):
[…]
4 return $\kappa$
5 else:
6 return $c$
Figure 4. A runtime shield synthesized by Aegis for the neural policy in a quadcopter system. Line 2 (marked in red) is a linear policy that outputs safe commands, while line 3 (marked in green) identifies potentially unsafe commands. Line 4 returns the safe command if the one from the neural policy is potentially unsafe.
As discussed in Section 1, the runtime shields created by VRL still have limitations regarding efficiency and permissiveness, as evidenced by their high runtime overhead and frequent unnecessary interventions in safe commands. Therefore, we introduce Aegis to improve the state-of-the-art. To give a sense of how our shields work compared to VRL’s, Figure 4 shows a shield synthesized by Aegis. The conditional statement in line 3 (marked in green) is a linear inequality, which offers improved computational efficiency over VRL’s barrier certificate with high-degree polynomials. On the other hand, the linear inequality measures how far away the neural policy output is from the output of the safe linear policy (line 2, marked in red), which has been formally verified. This direct comparison of the neural policy output to a safe output, rather than to an output of a polynomial barrier function that was fitted to bound the safe outputs, leads to an improved permissiveness of shields computed by our approach.
3. Aegis
This section provides details of Aegis, with its workflow illustrated in Figure 5. Aegis begins by constructing a program sketch that consists of a linear policy and a switching strategy (Section 3.1). Then, given a complex and possibly unknown physical environment, Aegis derives its state transition function by numerical differentiation (Griewank, 2013) (Section 3.2). Next, Aegis initializes a linear policy using the Linear-Quadratic Regulator (LQR) algorithm (Anderson and Moore, 2007) and iteratively refines it with a counterexample-guided inductive synthesis algorithm, until the linear policy is verified to be safe against the safety specifications (Section 3.3), determining the parameters of the linear policy. Following this, Aegis synthesizes the threshold of the switching strategy using a Bayesian optimization algorithm, aiming to minimize unnecessary shield interventions (Section 3.4). Finally, the synthesized components are incorporated into the program sketch to form a programmatic runtime shield. The following subsections describe each step in detail.
Figure 5. The workflow of Aegis.
3.1. Program Sketch
In Aegis, we model the synthesis of programmatic runtime shields as a sketch-based program synthesis problem (Solar-Lezama, 2009, 2008). A sketch is a partial code template with blank holes to be filled by synthesized parameters or statements. Considering that the synthesized shields should impose minimal overhead on the system, we set our program sketch to include only linear functions and inequalities, rather than following Zhu et al.’s work (Zhu et al., 2019) to use barrier certificates with high-order polynomials. This is because linear functions and inequalities are more efficient in terms of computational cost and are also widely used in control systems (Mania et al., 2018). Another advantage of using linear functions is that the computational cost of synthesizing them is relatively lower compared to obtaining barrier certificates by solving optimization problems with constraint solvers (Zhu et al., 2019; Nguyen et al., 2017).
We construct the program sketch as follows:
1def programmatic_runtime_shield($c$, $x_{1}$, $x_{2}$, …, $x_{n}$):
[…]
4 return $\mathbf{K}$
5 else:
6 return $c$
This sketch partially implements a runtime shield that takes the command $c$ from the neural policy and the environment state $x_{1},x_{2},...,x_{n}$ as input and returns a safe command. There is a linear policy that produces a command $\mathbf{K}$ based on the environment state variables (line 2, marked in green), where $\theta_{1},\dots,\theta_{n}$ are coefficients in the linear policy to be synthesized. Another main component is a conditional statement, which is a switching strategy that determines when to invoke the linear policy to replace the unsafe command $c$ with the safe command $\mathbf{K}$ (line 3, marked in red), where $\lambda$ is the threshold in the switching strategy to be synthesized. The sketch is then instantiated by filling the holes with the parameters synthesized by Aegis.
3.2. Inferring Environment Dynamics
In real-world applications, the environments in which neural policies operate are often complex and characterized by unknown dynamics functions (Levine and Abbeel, 2014), making it challenging or infeasible to directly verify the effectiveness of the synthesized programmatic runtime shields (Zhu et al., 2019). Therefore, Aegis infers the dynamics function of the physical environment in which the neural policy operates, making the unknown environment dynamics explicit and thereby facilitating the subsequent synthesis and verification of the linear policy.
As defined in Section 2, an environment state $s_{t}$ transitions to the subsequent state $s_{t+1}$ based on the state transition function $s_{t+1}=s_{t}+f(s_{t},c_{t})\times\Delta t$, where $c_{t}$ represents the commands issued by the neural policy at time step $t$. What we aim to infer is $f(s_{t},c_{t})$, which describes the dynamics of the physical environment, i.e., how the next state evolves in response to both the commands from the neural policy and the current state. We adopt the assumption that the environment dynamics can be approximated by a discrete-time linear time-invariant system model (Rugh, 1996; Zadeh and Desoer, 2008), which is widely used for modeling the dynamics of physical environments (Menghi et al., 2020; Zhu et al., 2019; Lillicrap et al., 2016). Such a model can be defined as follows:
(4) $f^{\prime}(s_{t},c_{t})=\mathbf{A}s_{t}+\mathbf{B}c_{t}\approx f(s_{t},c_{t})$
where $\mathbf{A}$ and $\mathbf{B}$ are two constant matrices. The linear time-invariant system model $f^{\prime}(s_{t},c_{t})$ is a linear approximation of the unknown dynamics function $f(s_{t},c_{t})$ around an equilibrium point (Cooke and Huang, 1996; Ramos, 2005), defined by specific environment states and commands $s^{*}$ and $c^{*}$ such that applying $c^{*}$ to $s^{*}$ results in no change in the environment state over time, i.e., $f(s^{*},c^{*})=0$. In linearization, the equilibrium point is often chosen as the initial environment state $s_{0}$ and the command $c_{0}$ at the beginning of the system operation, i.e., $s^{*}=s_{0}$ and $c^{*}=c_{0}$, so that the system’s behavior can be approximated by a linear model right from the start, simplifying the subsequent synthesis and verification processes. In such a case, the matrices $\mathbf{A}$ and $\mathbf{B}$ in $f^{\prime}(s_{t},c_{t})$ can be derived as the Jacobian matrices (Arrowsmith and Place, 1992), which represent the first-order partial derivatives of $f(s_{t},c_{t})$ with respect to $s_{t}$ and $c_{t}$, respectively. Considering $f(s_{t},c_{t})$ as a vector-valued function with $s_{t}$ and $c_{t}$ as input vectors having $m$ and $n$ dimensions and $f(s_{t},c_{t})$ as a $z$-dimensional output vector, respectively, then the Jacobian matrices $\mathbf{A}$ and $\mathbf{B}$ are $z\times m$ and $z\times n$ matrices, respectively, with the elements defined as:
(5) $\begin{gathered}\mathbf{A}_{ij}=\frac{\partial f_{i}\left(s_{t},c_{t}\right)}{\partial s_{j}},\quad\mathbf{B}_{ik}=\frac{\partial f_{i}\left(s_{t},c_{t}\right)}{\partial c_{k}}\\ i=1,\ldots,z,\quad j=1,\ldots,m,\quad k=1,\ldots,n\end{gathered}$
where $f_{i}(s_{t},c_{t})$ denotes the $i$-th element of the output vector $f(s_{t},c_{t})$, $s_{j}$ and $c_{k}$ are the $j$-th element of the state vector $s_{t}$ and the $k$-th element of the command vector $c_{t}$, respectively. The $\mathbf{A}_{ij}$ and $\mathbf{B}_{ik}$ are the elements of the Jacobian matrices $\mathbf{A}$ and $\mathbf{B}$, respectively.
Require: $f$, $s_{0}$, $c_{0}$, $z$, $m$, $n$
1
[…]
3 for $i\leftarrow 1$ to $z$ do
4       for $j\leftarrow 1$ to $m$ do
5             $\mathbf{A}_{ij}\leftarrow\big{(}f_{i}(s_{0}+\epsilon e_{j},c_{0})-f_{i}(s_{0}-\epsilon e_{j},c_{0})\big{)}/(2\times\epsilon)$
6       end for
7      for $k\leftarrow 1$ to $n$ do
8             $\mathbf{B}_{ik}\leftarrow\big{(}f_{i}(s_{0},c_{0}+\epsilon e_{k})-f_{i}(s_{0},c_{0}-\epsilon e_{k})\big{)}/(2\times\epsilon)$
9       end for
10
11 end for
return $\mathbf{A},\mathbf{B}$
Algorithm 1 Inferring Environment Dynamics
We obtain $\mathbf{A}$ and $\mathbf{B}$ using numerical differentiation (Gautschi, 2011; Cullum, 1971), which is commonly used to approximate a function’s derivative from a limited set of input-output pairs (Cheng et al., 2007). Specifically, it estimates the derivative by evaluating the function at two distinct inputs, computing the difference between the two outputs, dividing the difference by the distance between the two inputs, and then repeating the process for each input dimension. Formally, given the unknown dynamics function $f(s_{t},c_{t})$, the equilibrium point $s_{0}$ and $c_{0}$, the linearization algorithm computes the matrices $\mathbf{A}$ and $\mathbf{B}$ as follows:
(6) $\begin{gathered}\mathbf{A}_{ij}\approx\frac{f_{i}\left(s_{0}+\epsilon e_{j},c_{0}\right)-f_{i}\left(s_{0}-\epsilon e_{j},c_{0}\right)}{2\times\epsilon},\\ \mathbf{B}_{ik}\approx\frac{f_{i}\left(s_{0},c_{0}+\epsilon e_{k}\right)-f_{i}\left(s_{0},c_{0}-\epsilon e_{k}\right)}{2\times\epsilon}\\ i=1,\ldots,z,\quad j=1,\ldots,m,\quad k=1,\ldots,n\end{gathered}$
where $e_{j}$ and $e_{k}$ are unit vectors with the $j$-th and $k$-th elements being 1 and the rest being 0, respectively, and $\epsilon$ is a small random perturbation.
Algorithm 1 shows the pseudocode of how Aegis infers the linear approximation $f^{\prime}(s_{t},c_{t})$ of the unknown environment dynamics function via numerical differentiation. The algorithm takes as input the unknown dynamics function $f(s_{t},c_{t})$ (only its inputs and outputs can be accessed), the environment state $s_{0}$ and the command $c_{0}$ at the equilibrium point, the number of output dimensions $z$, the number of state dimensions $m$, and the number of command dimensions $n$. The algorithm first randomly produces a small perturbation $\epsilon$ (line 1) to slightly alter the environment state $s_{0}$ and the command $c_{0}$ around the equilibrium point. Subsequently, Equation 6 is applied to compute the elements of the Jacobian matrices $\mathbf{A}$ and $\mathbf{B}$ by evaluating the unknown dynamics function $f(s_{t},c_{t})$ at the perturbed inputs (lines 1-1). Finally, the algorithm returns the matrices $\mathbf{A}$ and $\mathbf{B}$ (line 1) to form the linear approximation $f^{\prime}(s_{t},c_{t})$ of the dynamics function $f(s_{t},c_{t})$.
3.3. Counterexample-Guided Inductive Synthesis
With the inferred environment dynamics function, we are now prepared to synthesize the linear policy, which is the first major component of the programmatic runtime shield we create. Our goal is to synthesize a linear policy that ensures all outputs adhere to the safety specifications. To achieve this, we introduce a counterexample-guided inductive synthesis algorithm, which is a popular approach for synthesizing programs from specifications (Češka et al., 2019; Solar-Lezama, 2009, 2008). This algorithm iteratively refines a candidate program using counterexamples generated during verification, ensuring that the synthesized program satisfies the desired properties. We tailor this algorithm to synthesize the linear policy for the programmatic runtime shield. Specifically, our algorithm initializes a linear policy using the Linear-Quadratic Regulator (LQR) algorithm (Anderson and Moore, 2007) and iteratively refines it based on counterexamples identified by the maximal output admissible set (Gilbert and Tan, 1991). We introduce our algorithm in the following paragraphs.
The LQR algorithm is a classic and effective method for designing optimal control policies for linear time-invariant systems (Anderson and Moore, 2007). It is widely used in various fields, including robotics (Mason et al., 2016), aerospace (Yang, 2014), and economics (Yaghobipour and Yarahmadi, 2020), due to its simplicity and effectiveness in managing multi-variable control systems. LQR derives control policies by minimizing a quadratic cost function, which is the sum of state and command costs. Specifically, given the state transition function $s_{t+1}=s_{t}+f(s_{t},c_{t})\times\Delta t$, where $f(s_{t},c_{t})=\mathbf{A}s_{t}+\mathbf{B}c_{t}$, the objective of the LQR is to determine the control input that minimizes the following quadratic cost function:
(7) $J=\sum_{t=0}^{\infty}\left(s_{t}^{T}\mathbf{Q}s_{t}+c_{t}^{T}\mathbf{R}c_{t}\right)$
where the term $s_{t}^{T}\mathbf{Q}s_{t}$ represents the state cost, and $c_{t}^{T}\mathbf{R}c_{t}$ represents the command cost, with $s_{t}$ and $c_{t}$ being the state and command at time step $t$, respectively. The intuition behind these costs is that the state cost penalizes deviations of the system’s current state from the desired state, encouraging the system to behave as expected, while the command cost penalizes the control commands, discouraging excessive actions and promoting smooth, efficient control that does not overly strain the system. The LQR algorithm computes the optimal control commands as a linear function of the current states, i.e.,
(8) $c_{t}=-\mathbf{K}s_{t}$
where $\mathbf{K}$ is the matrix of the parameters in the linear function, and is computed by solving the continuous-time Algebraic Riccati Equation (Anderson and Moore, 2007; Benner et al., 2020):
(9) $\mathbf{A}^{T}\mathbf{P}+\mathbf{PA}-\mathbf{PBR}^{-1}\mathbf{B}^{T}\mathbf{P}+\mathbf{Q}=0$
Once $\mathbf{A}^{P}$ is obtained, the matrix $\mathbf{K}$ can be computed as:
(10) $\mathbf{K}=\mathbf{R}^{-1}\mathbf{B}^{T}\mathbf{P}$
Note that in our case we assume that the state cost matrix $\mathbf{Q}$ and the command cost matrix $\mathbf{R}$ are well-defined and provided by the user, and how to choose them is beyond the scope of this paper. We refer interested readers to (Anderson and Moore, 2007) for more details. Moreover, in our work the above calculation is done with an off-the-shelf solver, i.e., pyMPC (Marcucci, 2024).
After initializing the linear policy using LQR, our counterexample-guided inductive synthesis algorithm iteratively refines the policy until it is verified to be safe according to the safety specifications. To facilitate this process, we use the maximal output admissible set (Gilbert and Tan, 1991) to identify counterexamples for further refinement. The maximal output admissible set is defined as the set of initial environment states for which the outputs of the linear policy are guaranteed to be safe commands. In other words, no matter which initial state within this set the environment starts from, the linear policy will never generate commands that could put the system in unsafe states at any time step. This set, denoted $\mathcal{X}$, can be defined as:
(11) $\mathcal{X}=\left\{s_{0}\in\mathcal{S}_{0}\mid\forall t\geq 0,s_{t}\in\neg\mathcal{S}_{u}\right.$
[…]
where $\mathbf{K}$ represents the parameters of the linear policy initialized by the LQR algorithm, $f^{\prime}$ is the inferred environment dynamics function obtained in Section 3.2, $\mathcal{S}_{0}$ is the set of initial environment states, and $t$ is the time step. Here, $\neg\mathcal{S}_{u}$ is used to indicate the complement of the set of unsafe states, meaning it encompasses the set of safe states that the system should remain in, thereby serving as the safety specifications for the system.
The maximal output admissible set is computed by solving linear programming problems (Gilbert and Tan, 1991). Specifically, considering the safety specifications $\lnot\mathcal{S}_{u}$ defined by a set of linear inequalities, i.e., $\lnot\mathcal{S}_{u}=\{s\mid D_{i}s\leq d_{i},i=1,\dots,n\}$, where $D_{i}$ is a matrix of coefficients for each state variable, $d_{i}$ is a vector of constants, and $n$ is the number of constraints, these inequalities define the boundaries of the set of safe states. Our goal is to ensure that the outputs of the linear policy remain within this set at every time step. Starting from the environment states $s_{1}$ at time step $t=1$, we check whether, given the dynamics function $f^{\prime}=\mathbf{A}s_{t}+\mathbf{B}c_{t}$ and the linear policy $c_{t}=-\mathbf{K}s_{t}$, the next state $s{2}=s_{1}+\mathbf{A}s_{1}-\mathbf{B}\mathbf{K}s_{1}$ remains within the set of safe states $\neg\mathcal{S}_{u}$ (note at $t=0$, verification is not needed since the next state is $s_{1}=s_{0}$, which is already within $\neg\mathcal{S}_{u}$ and thus safe by definition). For each constraint, the corresponding linear programming problem is as follows:
(12) $V(t=1,i)=\max_{s\in\neg\mathcal{S}_{u}}\left(D_{i}(s+\mathbf{A}s-\mathbf{B}\mathbf{K}s)-d_{i}\right)$
If for all $i$, $V(t=1,i)<0$, then the next state $s_{2}$ remains within the set of safe states $\neg\mathcal{S}_{u}$ for any state $s$ from $\neg\mathcal{S}_{u}$. According to the proved theorem in (Gilbert and Tan, 1991), this implies that all the future states of the system will also remain within the set of safe states $\neg\mathcal{S}_{u}$, provided with the initial state from $\neg\mathcal{S}_{u}$. Thus, we conclude the maximal output admissible set as $\mathcal{X}=\neg\mathcal{S}_{u}$. However, if one of the linear programming problems has $V(t=1,i)\geq 0$, it indicates that there exists $s_{1}$ and $c_{1}$ such that $s_{2}$ could go outside $\neg\mathcal{S}_{u}$. In such a case, we then proceed to the next time step and check if $s_{3}$ remains within $\neg\mathcal{S}_{u}$, and continue this process. If, at any time step, all values of $V(t,i)$ are less than 0, we can conclude that all the future states from time step $t$ will stay within the set of safe states $\neg\mathcal{S}_{u}$, with the updated maximal output admissible set as the intersection of $\neg\mathcal{S}_{u}$, $A\neg\mathcal{S}_{u}$, $A^{2}\neg\mathcal{S}_{u}$, …, up to the point where the iterative process stops. For ensuring $s_{t+1}$ stays within $\neg\mathcal{S}_{u}$, we need to check the following linear programming problem:
(13) $V(t,i)=\max_{\begin{subarray}{c}s\in\neg\mathcal{S}_{u}\cap A\neg\mathcal{S}_{u}\\ \cap\cdots\cap A^{t}\neg\mathcal{S}_{u}\end{subarray}}(D_{i}(s+\sum_{k=0}^{t-1}(\mathbf{A}s_{k}-\mathbf{B}\mathbf{K}s_{k})\cdot k)-d_{i})$
Although the above process sounds complex, our implementation leverages an off-the-shelf tool, i.e., pyMPC (Marcucci, 2024), which can efficiently solve the linear programming problems and compute the maximal output admissible sets. Note that in some extreme cases, there may be no time step $t\geq T$ where all $V(t,i)$ are less than 0, indicating a failure to compute the maximal output admissible set. In such scenarios, our method cannot derive runtime shields, leading to the termination of the synthesis process. However, in our experiments, this situation did not occur; the maximal output admissible set was successfully computed in all instances.
The utilization of the maximal output admissible set here is to identify counterexamples that show the linear policy is not universally safe. Intuitively, if the maximal output admissible set covers all initial states of the environment, the linear policy is theoretically verified to be safe, because as long as the system starts from an initial state within the maximal output admissible set, the linear policy will never generate unsafe commands at any time step, ensuring the system remains in safe states. However, if the maximal output admissible set does not cover all initial states, there exist counterexamples—specific initial states that cause the linear policy to produce unsafe commands at certain time steps. Formally, given the derived maximal output admissible set $\mathcal{X}$ and the set of initial environment states $\mathcal{S}_{0}$, the counterexamples can be defined as:
(14) $\mathcal{Z}=\left\{s_{0}\in\mathcal{S}_{0}\mid s_{0}\notin\mathcal{X}\right\}$
where $\mathcal{Z}$ is the set of counterexamples. If $\mathcal{Z}$ is not empty, it means that there exists an initial state $s_{0}$ in the environment such that the linear policy applied to $s_{0}$ may cause the system to enter unsafe states $\mathcal{S}_{u}$ at some future time steps, proving that the policy is not universally safe. Otherwise if $\mathcal{Z}=\emptyset$, this implies that the linear policy is verified to be safe for all possible initial states $s_{0}$ in $\mathcal{S}_{0}$.
When counterexamples are found, we proceed to refine the linear policy until it is verified to be safe without any counterexamples. To achieve this, we propose using a gradient-based optimization algorithm, with the objective of maximizing the rewards provided by the environment while running the linear policy with the counterexamples. The intuition behind this objective is that the environment’s reward function is typically designed to encourage the system to remain in safe states and penalize it for entering unsafe states. Therefore, by using the rewards as feedback, the linear policy is expected to increasingly generate safe commands after refinement. Additionally, this optimization algorithm differs from previous works (Zhu et al., 2019) that refine the linear policy to resemble the neural policy, because the neural policy is not guaranteed to be safe, and thus directly mimicking the neural policy may give inappropriate feedback for refining the linear policy, which could potentially lead the linear policy to adopt unsafe behaviors.
Require: $\mathbf{K}$, $\mathcal{Z}$, $f^{\prime}$, $r$, $lr$, $T$
1 for $ce\in\mathcal{Z}$ do
[…]
3       $s_{0}\leftarrow ce$
4       for time step $\text{time step }t\leftarrow 1$ to $T$ do
5             $s_{t}=s_{t-1}+f^{\prime}(s_{t-1},-\mathbf{K}s_{t-1})*t$
6             $c_{t}\leftarrow-\mathbf{K}(s_{t})$
7             $c^{+}_{t}\leftarrow-(\mathbf{K}+\epsilon)s_{t}$, $c^{-}_{t}\leftarrow-(\mathbf{K}-\epsilon)s_{t}$
8             $D_{+}\leftarrow-(c^{+}_{t}-c_{t})+r(s_{t},c^{+}_{t})$, $D_{-}\leftarrow-(c^{-}_{t}-c_{t})+r(s_{t},c^{-}_{t})$
9             $\Delta\leftarrow(D_{+}-D_{-})/(2\times\epsilon)$
10             $\mathbf{K}\leftarrow\mathbf{K}+lr\times\Delta$
11       end for
12
13 end for
return $\mathbf{K}$
Algorithm 2 Refine Linear Policy
We present the refinement algorithm in Algorithm 2. The algorithm takes as input the linear policy parameters $\mathbf{K}$, the counterexample set $\mathcal{Z}$, the inferred environment dynamics $f^{\prime}$, the environment’s reward function $r$, the learning rate $lr$, and the total number of time steps $T$. For each counterexample $ce$ in the set $\mathcal{Z}$, the algorithm first randomly produces a small perturbation $\epsilon$ (line 2) to slightly alter the parameters of the linear policy $\mathbf{K}$ in the subsequent process. We then initialize the environment with the counterexample $ce$ (line 2) and run the system with the linear policy to evaluate the outputs of the linear policy, perturbed linear policies, and the rewards obtained from the environment (lines 2-2). With these outputs and rewards, we calculate their differences as an estimate of the gradients $\Delta$ for updating the linear policy (lines 2-2). The iteration continues until a given number of time steps is reached, and returns the refined parameters of the linear policy $\mathbf{K}$ (line 2). The algorithm intuitively refines the linear policy by maximizing the rewards obtained from the environment, which are designed to encourage the system to remain in safe states and penalize it for entering unsafe states. The small perturbation $\epsilon$ is used to estimate the gradients of the rewards with respect to the parameters of the linear policy (Mania et al., 2018; Zhu et al., 2019), which are then used to update the linear policy in the direction that maximizes the rewards.
We are prepared to present the overall counterexample-guided inductive synthesis algorithm, as shown in Algorithm 3. The algorithm takes as input the inferred environment dynamics function $f^{\prime}$, the initial linear policy $\mathbf{K}$, the set of initial environment states $\mathcal{S}_{0}$, the safety specifications of the environment $\lnot\mathcal{S}_{u}$, the reward function $r$, the learning rate $lr$, and the total number of time steps $T$. It first computes the maximal output admissible set $\mathcal{X}$ (line 3) and finds the set of counterexamples $\mathcal{Z}$ (line 3). The algorithm then iteratively refines the linear policy using the counterexamples and updates the maximal output admissible set, until no more counterexamples are found and the linear policy is verified to be safe against the safety specifications (lines 3-3). The algorithm finally returns the synthesized parameters $\mathbf{K}$ for the linear policy (line 3).
Require: $f^{\prime}$, $\mathbf{K}$, $\mathcal{S}_{0}$, $\lnot\mathcal{S}_{u}$, $r$, $lr$, $T$
1 $\mathcal{X}\leftarrow\text{maximal\_output\_admissible\_set}(f^{\prime},\mathbf{K},\lnot\mathcal{S}_{u})$
2 $\mathcal{Z}\leftarrow\left\{s_{0}\in\mathcal{S}_{0}\mid s_{0}\notin\mathcal{X}\right\}$
3 while $\mathcal{Z}\neq\emptyset$ do
4       $\mathbf{K}\leftarrow\text{refine\_linear\_policy}(\mathbf{K},\mathcal{Z},f^{\prime},r,lr,T)$
5       $\mathcal{X}\leftarrow\text{maximal\_output\_admissible\_set}(f^{\prime},\mathbf{K},\lnot\mathcal{S}_{u})$
6       $\mathcal{Z}\leftarrow\left\{s_{0}\in\mathcal{S}_{0}\mid s_{0}\notin\mathcal{X}\right\}$
7
8 end while
9return $\mathbf{K}$
Algorithm 3 Counterexample-Guided Inductive Synthesis
3.4. Switching Strategy
The switching strategy is the second main component of a programmatic runtime shield that we synthesize. In the program sketch, the switching strategy is a conditional statement that determines when to replace the potentially unsafe command with the safe command produced by the linear policy. The linear policies are typically less powerful than the neural policies in terms of maintaining stable systems (discussed in section 5). Therefore, invoking the linear policy too frequently to replace the commands from the neural policy may have a negative impact on the stability of the system, necessitating a switching strategy that minimizes unnecessary interventions to maintain permissiveness.
As shown in our program sketch (Section 3.1), the switching strategy has a threshold $\lambda$ to be synthesized, which is used to compare the difference between the command produced by the neural policy and the linear policy. If the difference is greater than the threshold $\lambda$, the linear policy is invoked to replace the command from the neural policy. A smaller threshold $\lambda$ makes the switching strategy more sensitive to the difference, leading to more frequent interventions of the linear policy, while a larger threshold $\lambda$ makes the switching strategy too permissive and may fail to correct the unsafe commands from the neural policy. Thus, the threshold $\lambda$ should be carefully chosen to balance the trade-off, which is essentially an optimization problem.
We propose using a Bayesian optimization algorithm to solve this problem. Bayesian optimization is a powerful tool for optimizing black-box functions that are expensive to evaluate (Frazier, 2018; Snoek et al., 2012). It maintains a surrogate model, which approximates the real objective function but at a much lower computational cost. During the optimization process, candidate solutions are evaluated using the surrogate model, and the real objective function is evaluated only when a promising solution is identified. This approach significantly reduces the cost of evaluating candidate solutions, making it particularly well-suited for computationally expensive optimization problems.
In our case, we aim to optimize the threshold $\lambda$ of the switching strategy to minimize the number of violations of the safety specifications while maximizing the ratio of necessary interventions to total interventions. Formally, denoting the number of violations as $\mathcal{V}$, the number of necessary interventions as $\mathcal{V}^{*}$, and the total number of interventions as $\mathcal{I}$, the objective function can be formulated as:
(15) $\arg\min_{\lambda}\left(\log(\mathcal{V}+1)-\frac{\mathcal{V}^{*}}{\mathcal{I}+1}\right)$
where the $log$ function can make the optimization more sensitive to the small number of violations, ensuring that the algorithm focuses on minimizing the number of violations. The adding of 1 in the logarithm and the denominator is to avoid undefined values when the number of violations or interventions is zero. Intuitively, the ideal threshold $\lambda$ derived by the optimization algorithm with this objective function should minimize the number of violations to zero while maximizing the ratio of necessary interventions to total interventions, ensuring that the system remains safe with minimal interventions from the linear policy.
Note that this objective function is expensive to evaluate, because $\mathcal{V}$, $\mathcal{V}^{*}$, and $\mathcal{I}$ are unknown and cannot be directly computed using $\lambda$. Instead, they can only be obtained by running the environment with the linear policy and the neural policy to count the number of violations and interventions. Therefore, in our Bayesian optimization algorithm, we use the Gaussian process (GP) as a surrogate model of the objective function, which is a widely-used model in Bayesian optimization (Binois and Wycoff, 2022). The GP maintains a Gaussian distribution over the objective function to approximate the true function, which requires only a few evaluations of the objective function to guide the search for the optimal solution. Specifically, the GP can be modeled as:
(16) $obj(\lambda)\sim\mathcal{GP}(\mu(\lambda),k(\lambda,\lambda^{\prime}))$
where $obj(\lambda)$ is the objective function that we aim to optimize, $\mu(\lambda)$ and $k(\lambda,\lambda^{\prime})$ are called the mean function and the kernel function, respectively. We follow the implementation of the Gaussian process in the scikit-learn library (scikit optimize, 2024) to construct the model with the zero-mean function and the Matern kernel function (Genton, 2001). With each evaluation of the objective function, the GP model updates its parameters to better approximate the true function. Once the GP model is constructed and updated, it can predict the output of the objective function at a new input $\lambda$ as $\mu(\lambda)$. The Bayesian optimization algorithm iteratively refines the GP model and selects the next candidate solution to evaluate based on the predicted output of the GP, which guides the search for the optimal solution.
In addition, the acquisition function of the Gaussian process is used to select the next candidate solutions over the current best solution. We follow (Chen et al., 2021) to use expected improvement (Benassi et al., 2011) as the acquisition function, which measures the expected improvement of the surrogate model over the current best solution and selects the candidate solution that maximizes the expected improvement. Formally, the expected improvement is defined as:
(17) $\text{EI}(\lambda)=\mathbb{E}[\max(0,\mu(\lambda)^{*}-\mu(\lambda))]$
where $\mu(\lambda)^{*}$ is the current best solution, $\mu(\lambda)$ is the predicted output of the Gaussian process at input $\lambda$, and $\mathbb{E}$ is the expectation. With the acquisition function, in each iteration, the Bayesian optimization algorithm selects the candidate solution that maximizes the expected improvement, evaluates the objective function at the candidate solution, and updates the GP model to better approximate the true function. The iteration continues until the number of iterations is reached, and the algorithm returns the synthesized threshold $\lambda$.
4. Evaluation
This section presents the evaluation of Aegis. Our experiments aim to answer the following research questions:
  • RQ1 (Effectiveness): Can the shields synthesized by Aegis effectively maintain the safety of the neural policy?
  • RQ2 (Efficiency): What is the performance overhead of Aegis’s shields in terms of time and memory?
  • RQ3 (Permissiveness): How permissive are Aegis’s shields in terms of imposing shield interventions as few as possible?
  • RQ4 (Scalability): How scalable is the shield synthesis process of Aegis?
Table 1. Control systems and their safety specifications. The detailed description of these systems and specifications are in Section 4.1. The table also shows the number of inputs, outputs, and neural networks sizes of neural policies.
Control System Safety Specifications # Inputs/ # Outputs Neural Policy Size
#Layers #Neurons (M, a million)
Quadcopter $|\eta_{1}|<\pi/2\wedge|\eta_{2}|<\pi/2$ 2/1 2 0.4 M
CartPole $|\eta|<90^{\circ}\wedge|\delta|<0.3\wedge$ $|v_{1}|<0.3\wedge|v_{2}|<0.5$ 4/1 3 0.7 M
Self-Driving $|\eta|<90^{\circ}\wedge|d|<2.0$ 2/1 3 0.7 M
Pendulum $|\eta|<90^{\circ}\wedge|\omega|<\pi/2$ 2/1 3 7.0 M
4.1. Experimental Setup
Experimental Subjects
To comprehensively evaluate Aegis, we select four control systems that are heavily utilized in previous literature (Zhu et al., 2019), including autonomous driving navigation (Self-Driving), aerial drone control (Quadcopter) and dynamic equilibrium control (CartPole and Pendulum). Table 1 provides a brief description of each control system and its safety property.
The Quadcopter control system aims to realize a stable flight of a quadcopter, in which safety is ensured by limiting the angles of both the pitch and yaw rotations to not exceed $\pi/2$. The CartPole system consists of a pole attached to an unactuated joint connected to a cart moving along a frictionless track. The system is safe if the pole’s angle $\eta$ is no more than 90 from being upright, and the distance $\delta$ of the cart from the origin is less than 0.3 meters. Moreover, the velocity of the pole’s tip and cart should be less than 0.3 and 0.5 meters/s, respectively. The Pendulum system aims to balance a pendulum in an upright position. The safety is ensured by limiting the angle $\eta$ of the pendulum to not exceed 90 and the angular velocity $\omega$ to not exceed $\pi/2$. The Self-Driving system drives a car moving at a constant speed, and the safety property of this system is to ensure that the car never veers into canals found on both sides of the road, i.e, it is safe if the car’s heading angle $\eta$ remains within a threshold of $90^{\circ}$ and the distance $d$ between the car’s front and the road’s centerline stays below $2.0$ meters.
While these systems may appear simple, they are still representative of practical control systems that are difficult to control and maintain safety, such as Segway transporters (Segway, [n.d.]), camera drones (DJI, [n.d.]), and autonomous driving cars (Udacity, [n.d.]). Therefore, we believe that these control systems are appropriate for evaluating Aegis and the baseline.
We follow Zhu et al. (Zhu et al., 2019) to train the neural policies for the four control systems using the Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2016), a popular DRL algorithm for continuous control tasks. All hyperparameters are set to the default values from Zhu et al.’s official implementation. Note that our method is agnostic to the choice of DRL algorithms; other algorithms could also be applied. We choose DDPG because it is widely used in the literature (Wang et al., 2022) and aligns with the experimental setup of our baseline. We use the same neural network architecture as (Zhu et al., 2019) for each control system, which consists of 2 or 3 fully connected layers. The neural policies range in size from 0.4M to 7M neurons, as detailed in Table 1. Notably, these sizes are significantly larger than those in previous work, such as (Zhang et al., 2022; Tran et al., 2020), where policies typically contain no more than 400 neurons.
Experimental Settings and Metrics
For each control system, we evaluate the performance of Aegis and the baseline VRL (Zhu et al., 2019) by synthesizing programmatic runtime shields for the neural policies. Note that VRL’s shields apply barrier certificates with high-degree polynomials while our Aegis synthesizes shields with lightweight linear policies. We follow VRL’s original implementation to set the degree of the polynomials for the barrier certificates to 4, which is the value recommended by the authors. We also tried setting the degree to 1, i.e., making the barrier certificates linear, but VRL’s synthesis process failed to converge and cannot find shields that can guarantee safety. Therefore, we use the default setting of VRL to ensure a fair comparison.
We answer the research questions by comparing the performance of Aegis and baseline in terms of the following metrics. The effectiveness of each tool is measured by the number of safety violations found in the systems using the shields synthesized by each tool. The efficiency of the shields of each tool is measured by their time and memory overhead on the systems. The permissiveness of each tool is measured by the number of shield interventions required by each tool’s shields, as well as the ratio of the number of necessary shield interventions to the total number of interventions. The scalability of each tool is measured by the time required to synthesize a shield for the neural policy.
Additionally, to mitigate the impact of randomness, we repeat each experiment 10 times and report the average results along with their statistical significance and effect size as suggested by (Arcuri and Briand, 2011). We use the Mann-Whitney U test (Mann and Whitney, 1947) to measure the statistical significance and the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to measure the effect size. All the experiments are conducted on an Ubuntu 18.04 server equipped with an Intel Xeon E5-2698 CPU and 504 GB RAM.
4.2. RQ1: Effectiveness of Aegis
Table 2. Effectiveness of Aegis compared to the baseline. The table shows the number of safety violations found by Aegis and the baseline VRL.
System # Safety Violations
Neural Policy w/ VRL w/ Aegis
Quadcopter 182/1000 0/1000 0/1000
Pendulum 41/1000 0/1000 0/1000
Cart-Pole 47/1000 0/1000 0/1000
Self-Driving 35/1000 0/1000 0/1000
Average
76/1000
(7.6% Unsafe)
0/1000
(100% Safe)
0/1000
(100% Safe)
This RQ evaluates the effectiveness of Aegis in ensuring the safety of neural policies in the four control systems. We follow Zhu et al. (Zhu et al., 2019) to run the neural policy in each system over 1,000 runs, and each run is 5,000 time steps with random initial states. Then, we compare the number of safety violations found in the systems using neural policies, the policies with the baseline’s shields, and the policies with Aegis’s shields, as shown in Table 2.
We find that the neural policy violates the safety specifications in all the control systems, at a rate of 7.6% on average, i.e., 76 safety violations in 1,000 runs. However, after the Aegis’s and VRL’s shields are deployed, all the safety violations are eliminated, i.e., the number of violations in safety specifications is reduced to 0 in all the control systems, which suggests that both Aegis and the baseline can effectively enforce the safety by correcting all the unsafe commands and let the systems operate within the safety specifications.
RQ1 Takeaway.   Aegis can effectively ensure the safety of the neural policies in all control systems by correcting all the unsafe commands and letting the systems operate without any violations in safety specifications.
4.3. RQ2: Overhead of Aegis
Time Overhead
A programmatic runtime shield should be able to respond quickly without affecting other processes in the system, so we compare the time overhead of the shields synthesized by Aegis and VRL on the four control systems first. We measure the time overhead as the percentage of the additional operating time required by the deployed shields in the systems, i.e., if the original operating time of the system is 100s and it takes 101s after the shield is deployed, then the time overhead is 1% ((101-100)/100). Table 3 shows the time overhead of Aegis compared to the baseline. On average, the programmatic runtime shield synthesized by Aegis incurs only 5.0% time overhead, which is 5.5% lower than the 10.5% time overhead incurred by the baseline. The results show that Aegis improves the time overhead by 2.1$\times$ on average compared to the baseline.
To account for the impact of randomness and ensure the credibility of the results, we use the Mann-Whitney U test (Mann and Whitney, 1947), which is a non-parametric statistical test used to compare two independent groups. We use it to measure the statistical significance of the difference in time overhead between the runtime shields of the two tools across the 10 experiments. Our results show that Aegis significantly outperforms VRL on all four control systems at a 99% confidence level. We also use the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to measure the effect size. This non-parametric measure gives the probability that a randomly chosen value from one group is higher or lower than one from another group. It is commonly used to evaluate search-based testing methods due to their randomness (Palomba et al., 2016; Meng et al., 2022; Olsthoorn et al., 2020). According to Vargha and Delaney (Vargha and Delaney, 2000), $\hat{A}_{12}$ values greater than 0.71 (or less than 0.29) indicate a “large” effect size. Values between 0.64 and 0.71 (or between 0.29 and 0.36) correspond to a “medium” effect size, while values between 0.36 and 0.64 suggest a “negligible” effect size. As shown in Table 2, Aegis significantly outperforms VRL with a large effect size on four control systems.
Table 3. Overhead of Aegis compared to the baseline. The table shows the time and memory overhead of Aegis compared to the baseline. The last two columns show the statistical significance and effect size of the overhead.
Control System Time Overhead Memory Overhead
VRL Aegis p-value Â12 VRL Aegis p-value Â12
Quadcopter 12.1% 5.3% (2.3$\times$) 5.7e-3 < 0.01 L (0.97) 1.98 MB 0.53 MB (3.7$\times$) 0.03 < 0.05 M (0.73)
Pendulum 11.9% 4.1% (2.9$\times$) 9.1e-5 < 0.01 L (1.00) 4.05 MB 0.76 MB (5.3$\times$) 2.2e-4 < 0.01 L (0.97)
Cart-Pole 10.5% 4.9% (2.1$\times$) 8.5e-4 < 0.01 L (0.87) 1.49 MB 0.58 MB (2.6$\times$) 2.9e-4 < 0.01 L (0.96)
Self-Driving 7.4% 5.5% (1.3$\times$) 7.1e-3 < 0.01 L (0.80) 3.13 MB 0.57 MB (5.5$\times$) 6.0e-3 < 0.01 L (1.00)
Average 10.5% 5.0% (2.1$\times$) 2.66 MB 0.61 MB (4.4$\times$)
Memory Overhead
We then compare the memory overhead of the shields synthesized by Aegis and the baseline after deploying them into the four control systems. We measure the memory overhead as the percentage of the additional memory required by the deployed shields in the systems. Table 3 shows the memory overhead of Aegis compared to the baseline. On average, the programmatic runtime shield synthesized by Aegis occupies only 0.61 MB memory, while the baseline’s shields occupy 2.66 MB memory, which suggests that Aegis greatly reduces the memory overhead by 4.4 times compared to the baseline.
Similarly, we run each experiment 10 times and use the Mann-Whitney U test to measure the statistical significance and the Vargha-Delaney statistic $\hat{A}_{12}$ to measure the effect size of improvements. Our results show that Aegis significantly outperforms VRL on all four control systems at a 99% confidence level in terms of memory overhead, with a large or medium effect size.
RQ2 Takeaway.   The shields synthesized by Aegis significantly outperform those of the baseline in terms of time and memory overhead on all control systems. Specifically, Aegis’s shields incur only 5.0% time overhead and occupy only 0.61 MB memory on average, which are 2.1$\times$ and 4.4$\times$ improvements over the baseline, respectively.
4.4. RQ3: Permissiveness of Aegis
We first compare the permissiveness of the shields synthesized by Aegis and VRL in terms of the number of shield interventions over 1,000 runs (each run has 5,000 time steps) on the four control systems. Table 4 shows the number of shield interventions imposed by the shields of each tool. On average, Aegis requires only 384 shield interventions to ensure safety, which is 1.6$\times$ fewer than the 609 shield interventions required by the baseline. The statistical significance and effect size of the number of shield interventions are also shown in the last two columns of Table 4. Our results show that the improvements are statistically significant at the 99% confidence level with a large effect size on all four control systems.
In addition, we compare the ratio of the number of necessary shield interventions to the total number of interventions, shown in Figure 6. Here a necessary shield intervention is one that prevents a violation of the safety specifications, while an unnecessary intervention occurs when the safety specifications would not be violated even without the intervention. We observe that for each control system, Aegis’s shields impose a higher ratio of necessary shield interventions, suggesting that Aegis’s shields tend to intervene less frequently than the baseline’s shields and only when necessary. The ratio of necessary shield interventions to total number of interventions is on average 9.9% higher for Aegis than for the baseline. These results demonstrate that Aegis’s shields are more permissive than the baseline’s shields.
Table 4. Permissiveness of Aegis compared to the baseline. The table shows the number of shield interventions required by Aegis and the baseline. The last two columns show the statistical significance and effect size of the number of shield interventions.
System # Shield Interventions
VRL Aegis p-value Â12
Quadcopter 296 232 (1.3$\times$) 8.3e-3 < 0.01 L (0.92)
Pendulum 50 45 (1.1$\times$) 8.6e-5 < 0.01 L (1.00)
Cart-Pole 1941 1133 (1.7$\times$) 9.8e-5 < 0.01 L (1.00)
Self-Driving 147 127 (1.2$\times$) 6.1e-3 < 0.01 L (1.00)
Average 609 384 (1.6$\times$)
Figure 6. The ratio of necessary interventions to total number of interventions for Aegis and VRL.
RQ3 Takeaway.  Aegis’s shields significantly outperform those of baseline in terms of permissiveness on all control systems. Specifically, Aegis requires only 384 shield interventions on average, which is a 1.6$\times$ improvement over the baseline. Furthermore, Aegis’s shields have a 9.9% higher ratio of necessary shield interventions to total number of interventions than the baseline’s shields.
4.5. RQ4: Scalability of Aegis
In the last RQ, we evaluate the scalability of Aegis by comparing the time required to synthesize a shield for a neural policy. Table 5 shows the time cost of Aegis compared to the baseline. We observe that Aegis requires 65s on average, which reduces 77.3% of the time cost compared to the baseline. The statistical significance and effect size of the time cost also show that the improvements are statistically significant at the 99% confidence level with a large effect size on all 4 control systems.
Beyond Table 5, we also draw a line chart to compare the time required to synthesize a shield for neural policies of different sizes, shown in Figure 7. As Zhu et al. (Zhu et al., 2019) provides neural policies of different sizes for the four control systems, we use these neural policies to evaluate the scalability of Aegis. We divide the neural policies into three groups based on their sizes: small (0.4M neurons), medium (0.7M neurons), and large (7M neurons). Note that such sizes are sufficiently large for neural policies, especially when contrasted with policies used in previous work such as (Zhang et al., 2022; Tran et al., 2020), which typically contain no more than 400 neurons.
In Figure 7, we can see that the synthesis time of Aegis does not significantly increase with the size of the neural policies, while the baseline’s synthesis time increases greatly when the size of the neural policies increases. This result demonstrates that Aegis’s synthesis is more scalable than the baseline’s synthesis.
Table 5. Time cost of Aegis compared to VRL. The table shows the time required to synthesize a shield for the neural policies. The last column shows the statistical significance and effect size of the time cost.
System Time Cost of Shield Synthesis
VRL Aegis p-value Â12
Quadcopter 28s 14s (-50.0%) 2.2e-4 < 0.01 L (0.97)
Pendulum 439s 85s (-80.6%) 1.8e-4 < 0.01 L (1.00)
Cart-Pole 395s 118s (-70.1%) 9.3e-5 < 0.01 L (1.00)
Self-Driving 281s 43s (-84.7%) 9.5e-5 < 0.01 L (1.00)
Average 286s 65s (-77.3%)
Figure 7. Time required for the shield synthesis of neural policies at different sizes.
RQ4 Takeaway  Aegis consistently shows better scalability than the baseline on all control systems. Aegis requires 65s on average for synthesizing a shield for a neural policy, which reduces 77.3% of the time cost compared to the baseline. Furthermore, Aegis’s synthesis time does not significantly increase with the size of the neural policies.
5. Discussion
5.1. Runtime Shields vs. Neural Policies
One question that remains unanswered is:
Why do we need neural policies if runtime shields can control the systems?
To answer this question, we compare the performance of neural policies and runtime shields in terms of the number of steps required to reach steady states. Figure 8 shows the results.
The figure shows that neural policies consistently require fewer steps to reach steady states compared to our runtime shields, with an average of 10.5 steps for neural policies versus 19.3 steps for runtime shields. This suggests that neural policies are more efficient at maintaining system stability. Therefore, we conclude that neural policies are essential for control systems, while runtime shields serve primarily to ensure the safety of these policies. Additionally, we observe that neural policies paired with Aegis require more steps to reach steady states than those without shields, increasing the average steps from 10.5 to 15.75. This demonstrates that deploying runtime shields can compromise system stability, reflecting the “no free lunch” theorem (Wolpert and Macready, 1997), where safety comes at the cost of stability. This observation somewhat aligns with the finding of Zhu et al. (Zhu et al., 2019), which highlights that while alternative synthesis methods such as LQR-based approaches (Tedrake, 2010; Reist et al., 2016) can yield controllers with superior stability, they often fail to meet safety constraints. Our approach, therefore, represents a trade-off between safety and stability, but prioritizing safety. We leave the exploration of methods that can achieve better stability while maintaining safety to future work.
Figure 8. Time steps to reach steady state for the quadcopter system.
5.2. Threats to Validity
One threat to internal validity is the random behavior of the optimization algorithms used in our experiments. To mitigate this threat, we repeat each experiment 10 times and use appropriate statistical tests to account for both statistical significance and effect size, as recommended by Arcuri and Briand (Arcuri and Briand, 2011). Therefore, we believe that the randomness of the results is well-controlled. Another potential threat to internal validity is the implementation of Aegis and our scripts. To reduce this threat, we have carefully reviewed the code before submission and made it publicly available for review and improvement.
In terms of external validity, one potential threat is that the results of our analysis may not be fully generalizable. To mitigate this, we carefully selected a representative set of control systems with diverse tasks to evaluate our method and baseline under various conditions, ensuring that our results are not biased. However, it is important to note that our method relies on the assumption that the system dynamics are sufficiently linearizable when inferring environment dynamics (see Section 3.2), which may not be true for all control systems. Consequently, even though the control systems used in our experiments are non-linear, the generalizability of our method to a broader class of nonlinear systems warrants further investigation in future work.
One potential threat to construct validity is that the evaluation metrics may not fully capture the performance of falsification tools. To mitigate it, we used a total of six commonly-used evaluation metrics that are introduced in Section 4.1 to compare the performance of Aegis and the baseline from a comprehensive set of perspectives.
6. Related Work
This section reviews the related literature and positions our work in the context of existing research.
Runtime Shields for Neural Policies
Runtime shields have been used in runtime enforcement for reactive systems (Wu et al., 2019; Könighofer et al., 2017; Akametalu et al., 2014; Gillula and Tomlin, 2012; Bloem et al., 2015; Könighofer et al., 2020; Alshiekh et al., 2018) long before the advent of neural policies. However, runtime shields for neural policies are a relatively new concept and have only been studied in a few papers. Runtime shields in reinforcement learning models were first introduced in (Alshiekh et al., 2018). Because their verification approach is not symbolic, however, it can only work over finite discrete states instead of continuous state settings like the ones we consider in this paper. Zhu et al. (Zhu et al., 2019) proposed a method to synthesize deterministic and verified programs from neural policies, which can be used as runtime shields for neural policies. However, their shields have limitations in terms of efficiency and permissiveness, which necessitates the need for a more effective and permissive approach, motivating our work. Xiong et al. (Xiong and Jagannathan, 2021) proposed a method capable of synthesizing high-quality runtime shields even when the subject system involves hundreds of dimensions. However, their method is not effective enough to synthesize runtime shields that can guarantee safety for neural policies but still occasionally let unsafe commands pass through. Therefore, we do not consider it as a baseline in our experiments. Our approach differs from these studies in that we propose a new approach that can synthesize shields with lightweight overhead and high permissiveness, which is a novel and unique perspective in the context of runtime shields for neural policies.
Program Synthesis for Neural Policies
Although program synthesis has been studied for decades (Sobania et al., 2022), very little attention has been paid to program synthesis specialized for neural policies. A few studies (Verma et al., 2018; Trivedi et al., 2021) have proposed methods for generating programmatic policies from well-trained DRL models, which are more interpretable than neural networks. Additionally, Inala et al. (Inala et al., 2020) and Qiu et al. (Qiu and Zhu, 2021) synthesize programmatic controllers to perform continuous control tasks. Moreover, there are a few studies that focus on synthesizing code implementations or configurations for deep learning models in various tasks (Zhou et al., 2024; Shi et al., 2024; Liu et al., 2023; Shi et al., 2023). Our approach differs from these studies in that we propose a counterexample-guided inductive program synthesis approach which can efficiently synthesize runtime shields for neural policies, which is a unique application of program synthesis for neural policies.
Formal Verification of Neural Policies
Formal verification of neural policies has been explored in several studies (Landers and Doryab, 2023; Mazouni et al., 2023). Some of these studies align with our work by using counterexample-guided inductive synthesis (CEGIS) to formally verify neural policies and their neural certificates. Certain works, including (Chang et al., 2019; Abate et al., 2020, 2021; Edwards et al., 2024), focus on deterministic systems, employing symbolic reasoning to verify neural policies and construct Lyapunov functions (Haddad and Chellaboina, 2008) or barrier certificates in a CEGIS loop. Another more recent line of work (Lechner et al., 2022; Žikelić et al., 2023; Chatterjee et al., 2023; Ansaripour et al., 2023) considers both deterministic and stochastic systems, utilizing discretization and abstract interpretation for verification within the CEGIS loop, and deriving neural certificates of correctness or stability in the form of supermartingales instead of Lyapunov and barrier functions. Our work, however, differs by focusing on synthesizing runtime shields as simple programs for neural policies. This approach aims to enhance the efficiency and permissiveness of these shields, leading to a unique application of formal verification in the context of neural policies.
7. Conclusion and Future Work
In this paper, we present Aegis, a novel approach to synthesizing lightweight and permissive programmatic runtime shields for neural policies. Aegis achieves this by formulating the seeking of a runtime shield as a sketch-based program synthesis problem and using two novel algorithms to solve it, namely a counterexample-guided inductive synthesis algorithm and a Bayesian optimization algorithm. To evaluate Aegis and its synthesized shields, we use four representative control systems with neural policies and compare Aegis to the current state-of-the-art VRL. Our results show that the programmatic runtime shields synthesized by Aegis can correct all unsafe commands from neural policies, ensuring that the systems do not violate the desired safety properties at any time. In particular, compared to VRL, Aegis’s runtime shields exhibit a 2.1$\times$ reduction in time overhead and a 4.4$\times$ reduction in memory overhead, suggesting that they are much more lightweight. Additionally, Aegis’s shields show improved permissiveness, leading to an average of 1.6$\times$ fewer interventions compared to VRL’s shields. These results not only highlight the superiority of Aegis over the state-of-the-art, but also show that Aegis is an effective framework for synthesizing programmatic runtime shields for neural policies.
In the future, we plan to extend Aegis to handle more complex control systems such as those that take more complex inputs, including images and videos. We also plan to work on more advanced program synthesis techniques to further improve Aegis.
Replication Package.   For reproducibility and advancing the state of research, we have made our tool and all experimental data publicly available: https://github.com/jiekeshi/AEGIS.
Acknowledgements.
This research/project is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-017). We would like to thank the anonymous reviewers for their valuable feedback.